# Parametric MMD Estimation with Missing Values: Robustness to Missingness and Data Model Misspecification

### v1

@@ 6 Conclusion
In this paper, we adapted the concept of MMD estimation for missing values to obtain a parametric estimation procedure with provable robustness against misspecificiations of the data model as well as the missingness mechanism.
@@ Acknowledgements
Appendix A Proofs
A.1 Proofs from Section 3
@@ A.1.1 Intermediate lemmas
A.1.2 Consistency and asymptotic normality
@@
Invoking Theorem 2.1 in Newey and McFadden, (1994), Assumption 3.1 and 3.2 along with Lemma A.1 imply the strong consistency of the minimizer $\widehat{\theta}_{n}$ of $\widehat{L_{n}}$ towards the unique minimizer of $L$. ∎
@@
According to Lemma A.2 (3.4 and 3.5 are satisfied), the first term in the bound is going to zero almost surely, as well as the second one due to the continuity of ${\partial^{2}L(\cdot;)}/{\partial\theta_{k}\partial\theta_{j}}$ at $\theta_{\infty}^{\textnormal{MMD}}$ (3.4). Notice then that the definition of $B$ in 3.6 matches $\nabla_{\theta,\theta}^{2}L(\theta_{\infty}^{\textnormal{MMD}})$, and we obtain the almost sure convergence of the matrix $H_{n}\rightarrow B$ ($P^{*}$-a.s.) as $n\rightarrow+\infty$.
@@
A.2 Proofs from Section 4
@@
$\leq\frac{\mathbb{V}_{X\sim{\mathbb{P}_{X}}}\left[\pi_{m}(X)\right]}{\pi_{m}}.$
@@
We implicitly assume from now on that the quantity |$|\{i:M_{i}=0\}|$ in the denominator is actually equal to $\max(1,|\{i:M_{i}=0\}|)$ so that $P_{n}$ is well-defined. We then have for any $\theta\in\Theta$:
@@
[…]
@@
${\mathbb{E}}$ [${}_{M\sim{\mathbb{P}_{M}}}\left[\mathbb{D}^{2}\left({P^{(M)}_{\theta^{*}}},{\mathbb{P}^{(M)}_{X}}\right)\right]={\mathbb{E}}_{M\sim{\mathbb{P}_{M}}}\left[\mathbb{D}^{2}\left({P^{(M)}_{\theta^{*}}},(1-\epsilon){P^{(M)}_{\theta^{*}}}+\epsilon\mathbb{Q}_{X}^{(M)}\right)\right]$
@@
${\mathbb{E}}$ [${}_{M\sim{\mathbb{P}_{M}}}\left[\mathbb{D}^{2}\left(P_{\theta_{\infty}^{\textnormal{MMD}}}^{(M)},{P^{(M)}_{\theta^{*}}}\right)\right]\leq 6\cdot{\mathbb{E}}_{M\sim{\mathbb{P}_{M}}}\left[\mathbb{D}^{2}\left({P^{(M)}_{\theta^{*}}},{\mathbb{P}^{(M)}_{X}}\right)\right]+8\cdot{\mathbb{E}}_{M\sim{\mathbb{P}_{M}}}\left[\frac{\mathbb{M}_{X\sim{\mathbb{P}_{X}}}\left[\pi_{M}(X)\right]}{\mathbb{E}_{X\sim{\mathbb{P}_{X}}}\left[\pi_{M}(X)\right]^{2}}\right]$
@@
A.3 Proofs from Section 5
@@
With the properties ||$\big{|}|\{i:\widetilde{M}_{i}=0\}|-|\{i:M_{i}=0\}|\big{|}\leq\varepsilon/\alpha\cdot|\{i:M_{i}=0\}|$, |$|\{i:\widetilde{M}_{i}=0\}|\geq(1-\varepsilon/\alpha)\cdot|\{i:M_{i}=0\}|$, |$|\{i:M_{i}=0,\widetilde{M}_{i}=1\}|\leq\varepsilon/\alpha\cdot|\{i:M_{i}=0\}|$ and |$|\{i:M_{i}=1,\widetilde{M}_{i}=0\}|\leq\varepsilon/\alpha\cdot|\{i:M_{i}=0\}|$, we have:
@@
=|$=\frac{|\{i:M_{i}=0,\widetilde{M}_{i}=1\}|}{|\{i:M_{i}=0\}|}+\frac{|\{i:M_{i}=1,\widetilde{M}_{i}=0\}|}{|\{i:\widetilde{M}_{i}=0\}|}$
+|$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{|\{i:M_{i}=\widetilde{M}_{i}=0\}|\cdot\left||\{i:\widetilde{M}_{i}=0\}|-|\{i:M_{i}=0\}|\right|}{|\{i:M_{i}=0\}|\cdot|\{i:\widetilde{M}_{i}=0\}|}$
@@
=|$=\frac{|\{i:M_{i}=0,\widetilde{M}_{i}=1\}|}{|\{i:M_{i}=0\}|}+\frac{|\{i:M_{i}=1,\widetilde{M}_{i}=0\}|}{|\{i:\widetilde{M}_{i}=0\}|}+\frac{|\{i:M_{i}=\widetilde{M}_{i}=0\}|\cdot\varepsilon/\alpha}{\{i:\widetilde{M}_{i}=0\}|}$

### v2

@@ 7 Conclusion
In this paper, we adapted the concept of MMD estimation for missing values to obtain a parametric estimation procedure with provable robustness against misspecificiation of the data model as well as the missingness mechanism.
@@ Acknowledgements
Appendix A Additional Empirical Results
A.1 Gaussian Mean Estimation under MCAR missingness and Model Contamination
We consider here the same setting as in the main text, but with $\varepsilon=0$, i.e. the MCAR mechanism is not contaminated, such that ${\mathbb{P}}[M=m_{j}\mid X=x]=\alpha_{m_{j}}=0.25$ for $j\in\{1,\ldots,4\}$. Again this leaves an expected fraction of 1/4 points fully observed. Results are shown in Table 2. Even without MNAR contamination, it appears that the effect of outliers on the mean becomes worse with MCAR data; while in Table 1 of Chérief-Abdellatif and Alquier, (2022) the RMSE of the mean is around 2 for the case of $N(\mathbf{10},I_{d})$ and $\delta_{\mathbf{10}}$, it is more than 3 times higher in Table 2. Here, the first row corresponds to no contamination ($\epsilon=\varepsilon=0$). As with MNAR contamination, the MMD estimator is very competitive, only performing worse than the other robust methods in the notorious case of a $\delta_{\mathbf{1}}$ contamination. Notable, the RMSE values of the MMD in Table 2 are very close to the ones in Table 1, indicating the the MNAR contamination makes very little difference to the MMD.
Table 2: RMSE under contamination of the data distribution ($\epsilon=0.2$, $\varepsilon=0$). Estimated standard deviations are given in brackets.
$N(\mathbf{0},I_{d})$ $N(\mathbf{1},I_{d})$ $N(\mathbf{10},I_{d})$ $\delta_{\mathbf{1}}$ $\delta_{\mathbf{10}}$
MMD 0.21 0.57 0.25 0.84 0.24
(0.05) (0.09) (0.06) (0.12) (0.05)
MLE 0.16 0.66 6.34 0.66 6.27
(0.04) (0.09) (0.54) (0.08) (0.71)
Median 0.2 0.63 1.05 1.04 1.04
(0.04) (0.1) (0.13) (0.15) (0.15)
evfiltering 0.28 0.7 0.33 0.67 0.33
(0.06) (0.14) (0.08) (0.13) (0.06)
QUE 0.28 0.7 1.97 0.67 1.95
(0.06) (0.14) (6.28) (0.13) (6.15)
A.2 Increased Fraction of Contamination
Table 3 presents the same results as in Section 6, but with an increased contamination fraction $(\epsilon=\varepsilon=0.3)$. Similarly, Table 4 shows the results under MCAR with an increased data contamination ($\epsilon=0.3$, $\varepsilon=0$).
Table 3: RMSE under contamination of both the data distribution and missingness mechanism ($\varepsilon=\epsilon=0.3$). Estimated standard deviations are given in brackets.
$N(\mathbf{0},I_{d})$ $N(\mathbf{1},I_{d})$ $N(\mathbf{10},I_{d})$ $\delta_{\mathbf{1}}$ $\delta_{\mathbf{10}}$
MMD 0.21 0.86 0.3 1.28 0.29
(0.05) (0.11) (0.07) (0.11) (0.08)
MLE 0.16 0.99 9.92 1 9.75
(0.03) (0.08) (0.62) (0.08) (0.64)
Median 0.2 0.95 1.96 1.93 1.94
(0.05) (0.09) (0.22) (0.21) (0.23)
evfiltering 0.32 1.24 0.49 1.31 0.55
(0.07) (0.16) (0.17) (0.17) (0.17)
QUE 0.32 1.24 11.54 1.31 13.91
(0.07) (0.16) (14.38) (0.17) (14.63)
Table 4: RMSE under contamination of the data distribution ($\epsilon=0.3$, $\varepsilon=0$). Estimated standard deviations are given in brackets.
$N(\mathbf{0},I_{d})$ $N(\mathbf{1},I_{d})$ $N(\mathbf{10},I_{d})$ $\delta_{\mathbf{1}}$ $\delta_{\mathbf{10}}$
MMD 0.21 0.88 0.3 1.3 0.29
2 (0.04) (0.1) (0.06) (0.12) (0.06)
mean 0.16 0.96 9.4 0.97 9.57
4 (0.03) (0.08) (0.76) (0.08) (0.73)
median 0.21 0.92 1.79 1.85 1.87
6 (0.05) (0.09) (0.25) (0.23) (0.23)
evpruning 0.28 0.97 0.38 1 0.37
8 (0.07) (0.12) (0.11) (0.14) (0.09)
que 0.28 0.97 7.53 1 8.11
10 (0.07) (0.12) (12.35) (0.14) (12.64)
Appendix B Proofs
B.1 Proofs from Section 3
@@ B.1.1 Intermediate lemmas
B.1.2 Consistency and asymptotic normality
@@
Invoking Theorem 2.1 in Newey and McFadden, (1994), Assumption 3.1 and 3.2 along with Lemma B.1 imply the strong consistency of the minimizer $\widehat{\theta}_{n}$ of $\widehat{L_{n}}$ towards the unique minimizer of $L$. ∎
@@
According to Lemma B.2 (3.4 and 3.5 are satisfied), the first term in the bound is going to zero almost surely, as well as the second one due to the continuity of ${\partial^{2}L(\cdot;)}/{\partial\theta_{k}\partial\theta_{j}}$ at $\theta_{\infty}^{\textnormal{MMD}}$ (3.4). Notice then that the definition of $B$ in 3.6 matches $\nabla_{\theta,\theta}^{2}L(\theta_{\infty}^{\textnormal{MMD}})$, and we obtain the almost sure convergence of the matrix $H_{n}\rightarrow B$ ($P^{*}$-a.s.) as $n\rightarrow+\infty$.
@@
B.2 Proofs from Section 4
@@
$\leq\frac{\sqrt{\mathbb{V}_{X\sim{\mathbb{P}_{X}}}\left[\pi_{m}(X)\right]}}{\pi_{m}}.$
@@
We implicitly assume from now on that the quantity $|\{i:M_{i}=0\}|$ in the denominator is actually equal to $\max(1,|\{i:M_{i}=0\}|)$ so that $P_{n}$ is well-defined. We then have for any $\theta\in\Theta$:
@@
where the first inequality comes from the standard inequality $\textnormal{Variance}\leq\textnormal{Second order moment}$, the second one from the boundedness of the kernel, and the last one as a standard bound on the negative moments of a binomial random variable (Chao and Strawderman, , 1972) (${\mathbb{E}}[1/(1+\textnormal{Bin(n,p)})]\leq 1/np$). Note that we also used the inequality $1/x\leq 2/(1+x)$.
@@
${\mathbb{E}}$ $_{M\sim{\mathbb{P}_{M}}}\left[\mathbb{D}^{2}\left({P^{(M)}_{\theta^{*}}},{\mathbb{P}^{(M)}_{X}}\right)\right]={\mathbb{E}}_{M\sim{\mathbb{P}_{M}}}\left[\mathbb{D}^{2}\left({P^{(M)}_{\theta^{*}}},(1-\epsilon){P^{(M)}_{\theta^{*}}}+\epsilon\mathbb{Q}_{X}^{(M)}\right)\right]$
@@
${\mathbb{E}}$ $_{M\sim{\mathbb{P}_{M}}}\left[\mathbb{D}^{2}\left(P_{\theta_{\infty}^{\textnormal{MMD}}}^{(M)},{P^{(M)}_{\theta^{*}}}\right)\right]\leq 6\cdot{\mathbb{E}}_{M\sim{\mathbb{P}_{M}}}\left[\mathbb{D}^{2}\left({P^{(M)}_{\theta^{*}}},{\mathbb{P}^{(M)}_{X}}\right)\right]+8\cdot{\mathbb{E}}_{M\sim{\mathbb{P}_{M}}}\left[\frac{\mathbb{M}_{X\sim{\mathbb{P}_{X}}}\left[\pi_{M}(X)\right]}{\mathbb{E}_{X\sim{\mathbb{P}_{X}}}\left[\pi_{M}(X)\right]^{2}}\right]$
@@
B.3 Proofs from Section 5
@@
With the properties $\big||\{i:\widetilde{M}_{i}=0\}|-|\{i:M_{i}=0\}|\big|\leq\varepsilon/\alpha\cdot|\{i:M_{i}=0\}|$, $|\{i:\widetilde{M}_{i}=0\}|\geq(1-\varepsilon/\alpha)\cdot|\{i:M_{i}=0\}|$, $|\{i:M_{i}=0,\widetilde{M}_{i}=1\}|\leq\varepsilon/\alpha\cdot|\{i:M_{i}=0\}|$ and $|\{i:M_{i}=1,\widetilde{M}_{i}=0\}|\leq\varepsilon/\alpha\cdot|\{i:M_{i}=0\}|$, we have:
@@
$=\frac{|\{i:M_{i}=0,\widetilde{M}_{i}=1\}|}{|\{i:M_{i}=0\}|}+\frac{|\{i:M_{i}=1,\widetilde{M}_{i}=0\}|}{|\{i:\widetilde{M}_{i}=0\}|}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{|\{i:M_{i}=\widetilde{M}_{i}=0\}|\cdot\left||\{i:\widetilde{M}_{i}=0\}|-|\{i:M_{i}=0\}|\right|}{|\{i:M_{i}=0\}|\cdot|\{i:\widetilde{M}_{i}=0\}|}$
@@
$=\frac{|\{i:M_{i}=0,\widetilde{M}_{i}=1\}|}{|\{i:M_{i}=0\}|}+\frac{|\{i:M_{i}=1,\widetilde{M}_{i}=0\}|}{|\{i:\widetilde{M}_{i}=0\}|}+\frac{|\{i:M_{i}=\widetilde{M}_{i}=0\}|\cdot\varepsilon/\alpha}{\{i:\widetilde{M}_{i}=0\}|}$